# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ Abstract
This review examines data from hadron-nucleus collisions, primarily focusing on hard processes like Drell-Yan, heavy flavor and quarkonium production. It highlights observed modifications of particle yields as functions of momentum and rapidity, aiming to clarify the underlying QCD effects in cold nuclear matter (CNM). The paper outlines strategies for future experiments, including the Electron-Ion Collider (EIC), to distinguish between these effects. Key questions address the universality of suppression mechanisms and the role of non-perturbative physics, providing a road map for upcoming nuclear data.
@@ I Introduction
The root cause of this unsatisfactory situation lies in the multitude of nuclear effects that can influence the nuclear cross sections. The well known EMC effect [1] provided evidence that the nuclear parton distributions functions (nPDFs) in bound nucleons differ from those in free nucleons, modifying the effective parton densities probed in the collision [2, 3, 4, 5, 6, 7, 8, 9]. Most of the hadron-nucleus experiments primarily probe the leading-twist shadowing region of the target nucleus at low momentum fraction $x$, leading to a reduction of the particle yields relative to an incoherent sum over all the nucleons.
@@
Another effect, relevant at high center-of-mass energies, is due to the gluon saturation. Saturation prevents an unlimited increase of the gluon density at low $x$ values, which would otherwise violate perturbative unitarity. It has an important impact on particle production, resulting in additional suppression at very forward rapidities. This suppression is then similar to the suppression resulting from the purely non-perturbative leading-twist shadowing. Although saturation and shadowing originate from two different QCD mechanisms, differentiating between them remains a difficult task.
@@
In addition, the incoming or outgoing parton may lose some of its energy by a medium-induced gluon radiation when traversing the nucleus, known as Landau-Pomeranchuk-Migdal (LPM) energy loss  [10, 11]. Similarly, the incoming and outgoing partons could undergo energy loss, which is summed coherently and is known as Fully Coherent Energy Loss (FCEL)  [12, 13]. A related, but distinct, phenomenon arises from multiple soft scatterings of the parton traversing the nuclear medium. Each consecutive scattering contributes to a slight modification of the transverse momentum of the parton as well as the broadening of the particle production yields as a function of transverse momentum [14, 15, 16, 17, 18, 19]. Since the broadening scales with the number of scatterings, it typically grows linearly with the nuclear size.
@@
Furthermore, heavy mesons produced in the final state are subject to inelastic interactions with the nuclear medium. When their formation length is shorter than the size of the nucleus, the newly formed mesons may be dissociated via scattering with partons in a high-density environment [20, 21]. In addition, both the initially produced quarks pair and the nascent heavy meson can interact with the surrounding nucleons. The corresponding absorption cross section can vary significantly depending on the parameters of the interaction. Moreover, the quarkonium state may undergo further interactions with comoving hadrons or partons, referred to comovers [22, 23]. These interactions, which result in the break-up of the bound state, are affected by the local medium density. The effects of collisional dissociation, nuclear absorption, and comover interactions are strongly dependent on the length traversed through the nuclear medium and all lead to an attenuation in the observed heavy meson yield. In a typical high-energy hadron-nucleus experiment, several CNM effects may be present simultaneously. Their relative importance, however, depends strongly on the process considered, the kinematic regime, and the specific observables under scrutiny. The difficulties in quantitatively establishing the relative contributions of these effects add to the challenge of forming a coherent theoretical framework.
@@
The upcoming Electron-Ion Collider (EIC) [24] and future fixed-target experiments [25, 26] will offer unprecedented opportunities to probe CNM with greater precision and across a broader kinematic range. Leveraging these new experimental capabilities, we sek to narrow the gap between theoretical predictions and experimental observations, paving the way for a unified description of cold nuclear matter effects.
We will identify three key questions that future data should address:
@@
This paper is organized in eight sections. After a short Introduction in Sec. I, the most relevant fixed-target and collider data are reviewed in Sec. II. Section III discusses the present status of the nPDFs and gluon saturation at small $x$. The radiative energy loss effects are reviewed in Sec. IV. Final-state interactions including collisional dissociation, absorption, and comovers are discussed in Sec. V. Additional effects due to impact-parameter dependence and intrinsic charm are presented in Sec VI. Section VII explores the experimental opportunities opened by the newest physics facilities and particularly by the upcoming Electron-Ion Collider. A short conclusion is drawn in Sec. VIII.
@@ II Drell-Yan and heavy flavor data on nuclear targets
Hadron-nucleus collisions serve as a powerful tool to probe the properties of cold nuclear matter, offering unique insights into the dynamics of partonic interactions in the nuclear medium. Assuming the validity of collinear factorization, they provide a useful context for extracting and studying nPDFs. Processes such as Drell-Yan (DY) and hadron production are particularly valuable probes of these effects. Fixed-target experiments, which benefit from a large experimental rapidity, $y$, coverage due to the beam boost, primarily explore nuclear effects at higher values of the parton momentum fraction $x$, $x>10^{-3}$. In contrast, data from the Large Hadron Collider (LHC) extend these studies into the low-$x$ regime, even though they cannnot cover the full rapidity range available reaching $x$ values as small as $\sim 10^{-5}$, where gluon saturation is expected to become significant. The nuclear modification factor, $R_{hA}$, quantifies the impact of the nuclear medium on particle production by comparing cross sections in hadron-nucleus, $h+A$, to hadron-proton, $h+p$, collisions. Studied as a function of rapidity or Feynman $x$ ($x_{F}$) and transverse momentum, $p_{T}$, $R_{hA}$ provides insight into the evolution of nuclear effects across kinematic regimes.
Despite extensive measurements at fixed-target facilities and colliders, the mechanisms driving the nuclear effects remain ill-determined. It does not appear possible to precisely determine the contributions from all proposed cold nuclear matter effects based on only one type of measurements. Gaining a more quantitative insight into these effects requires a combined analysis of various processes. The current status of the main processes used to investigate cold nuclear matter effects—namely Drell-Yan, quarkonium and heavy-flavor production in hadronic collisions, as well as particle production in ultraperipheral collisions—is summarized below.
@@ II.1 Drell-Yan Data
At leading order (LO), the electromagnetic Drell-Yan (DY) process describes the production of a virtual photon, $\gamma^{\star}$, via quark-antiquark annihilation and its decay into a lepton-antilepton pair, $\gamma^{\star}\to\ell^{+}\ell^{-}$. Along with deep-inelastic scattering (DIS), this process provides complementary information about the partonic structure of both the beam particle and the target proton or nucleus. The DY cross section is well understood and can be calculated using fixed-order perturbative methods [27, 28, 29] or employing advanced approaches such as Collins-Soper-Sterman (CSS) resummation and transverse-momentum-dependent (TMD) factorization [30], which accurately describes the low $p_{T}$ portion of the dilepton spectrum. The nuclear dependence of the DY process is quantified by the ratio $R_{hA}^{\text{DY}}$, which compares the yields in $h+A$ collisions to those in $h+p$ or, occasionally, in $h+{\rm d}$ collisions.
@@
Most of the DY data come from fixed-target experiments primarily performed at CERN and Fermilab, spanning the center-of-mass energy range $15\leq\sqrt{s}\leq 38.7$ GeV. The earliest experiments [31, 32, 33], which used proton and pion beams at 225 GeV and measured only total cross sections, showed no evidence of nuclear effects. This conclusion was based on comparisons of $p+p$ and $p+A$ cross sections using the relation $\sigma_{pA}=A^{\alpha}\sigma_{pp}$, where the exponent $\alpha$ was found to be consistent with unity. More detailed comparisons between heavy and light targets were performed in the beginning of the 80s using a pion beam at CERN [34] and a proton beam at Fermilab [35]. In spite of the relatively large statistics, both experiments reported the absence of nuclear effects. A clear variation of $R_{hA}^{\text{DY}}$ from unity was observed [36] a few years later by the NA10 experiment at CERN, when comparing the data collected on hydrogen and tungsten targets. [^1] The nuclear modification ratio $R_{hA}$ can be related to $\alpha$ on the cross section level by $R_{hA}=\sigma_{pA}/(A\sigma_{pp})=A^{\alpha-1}$. The measured $R_{hA}^{\rm DY}$ as a function of the target parton momentum fraction, $x_{2}$, was found to agree with the recently discovered EMC effect [1]. A pronounced increase of $R_{hA}^{\text{DY}}$ around $p_{T}=2$ GeV (the “Cronin peak” [37]) was also observed [38]. The NA10 results were soon qualitatively confirmed by extensive measurements by the E772 experiment at Fermilab [39]. Data taken with a 800 GeV proton beam, the experiment compared results from a deuterium target with four other nuclear targets (C, Ca, Fe and W) in the $x_{F}$ region between 0.1 and 0.3.222Note that Feynman $x$, $x_{F}$, can be defined as $x_{F}=x_{1}-x_{2}$ the difference between the projectile momentum fraction, $x_{1}$ and the target momentum fraction, $x_{2}$, with a range $-1\leq x_{F}\leq 1$. Another definition is $x_{F}=p_{z}/p_{z}^{\text{max}}$, where $p_{z}$ is the momentum of the dilepton along the beam direction, and $p_{z}^{\text{max}}=\sqrt{s}/2$. The data were found to be consistent with unity, except at low $x_{2}$, where the suppression increased with the mass of the target, $A$, an effect fairly compatible with the nuclear leading-twist shadowing measured in DIS. An increase of $R_{hA}^{\text{DY}}$ for $p_{T}\gtrsim 1$ GeV with $A$ was also clearly observed. A much larger $x_{F}$ region, up to $x_{F}=0.95$, was covered by the E866 collaboration [40] using three targets: Be, Fe and W. The data thus extend to larger values of $x_{1}$ and to smaller values of $x_{2}$, providing increased sensitivity to nuclear effects in these two domains. At low $x_{2}$, the experimental Fe/Be and W/Be ratios are compatible with nuclear leading-twist shadowing as observed in DIS while for $x_{1}\gtrsim 0.7$ the observed suppression provided the first evidence for an incoming parton energy loss in the DY process [41, 42, 4].
@@
A large amount of DY data employing a 450 GeV proton beam was also collected by the NA50 collaboration at CERN [43]. When corrected for nuclear isospin, the resulting cross sections for a number of nuclei from Be to W perfectly scale with the number of nucleon-nucleon collisions, in line with previous data on proton and deuteron targets. More recently, preliminary data [44] from the E906 experiment at Fermilab (SeaQuest) using a proton beam at 120 GeV became available. The measured nuclear modification factor $R_{pA}^{\text{DY}}$ was found to be smaller than unity as a function of $x_{F}$, with a stronger suppression at large $x_{F}\sim 0.8$. The modification appears to be more pronounced for the W/C ratio than for Fe/C.
@@
Measurements of the dilepton continuum have been extended by several collider experiments. The PHENIX collaboration at RHIC [45] reported data for $p+p$ and $p+\mathrm{Au}$ collisions at $\sqrt{s_{NN}}=200$ GeV focusing on the dilepton mass region below 10 GeV and $p_{T}\lesssim M$. A slight enhancement in the $p_{T}$ spectra was observed at forward rapidity around $p_{T}=2$ GeV. Due to the low statistics, these data do not exhibit any clear nuclear effects.
@@
The study of the Drell-Yan continuum is complemented by measurements of $Z$ and $W^{\pm}$ vector boson production made by the LHC experiments, see recent results in $p+{\rm Pb}$ collisions by CMS [47, 48, 49, 46], ALICE [50, 51] and LHCb [52, 53] at 5.02 and 8.16 TeV. Within uncertainties, the ALICE data show only a small difference between the ${p+p}$ and $p+A$ cross sections. In contrast, the LHCb data indicate a slight suppression of the nuclear cross section in the forward region and an enhancement in the backward region compared to data in ${p+p}$ collisions. Both datasets are compatible with the nPDF calculations. This difference may partly arise from the distinct rapidity coverages: ALICE data probe $-4.46<y<-2.96$ and $3.532.03<y<$ 3.53, while LHCb data access slightly different rapidity regions with $-4.0<y<-2.5$ and $4.01.5<y<$ 4.0. The different kinematic ranges could influence the observed nuclear effects.
The Drell-Yan and $Z$, $W^{\pm}$ vector bosons data from the fixed-target and collider experiments discussed above are summarized in Table 1.
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Process Target ($A$) Dependence Ref.
@@
Table 1: List of data sets for Drell-Yan and $Z$, $W^{\pm}$ vector bosons production on nuclear targets.
@@ II.2 Quarkonium Data
Quarkonium production involves the formation of a heavy quark-antiquark bound state through hadronization, $h+A\to[Q\bar{Q}]+X$. Several approaches have been used to describe this process, including the Color Evaporation Model (CEM) [57, 58, 59, 60], the Improved Color Evaporation Model (ICEM) [61], the Color Singlet Model (CSM) [62], and Non-Relativistic Quantum Chromodynamics (NRQCD) [63]. The CEM and CSM have not be rigorously applied beyond hadroproduction. The ICEM has been successfully been applied to production and polarization of $J/\psi$ in $p+p$ and $e+p$ collisions [64, 65, 66, 67, 68, 69]. Work is ongoing to apply it to other quarkonium production processes. Although theoretically sound and phenomenologically successful, the NRQCD approach appears incomplete, as outlined in the next paragraph.
In the NRQCD framework, the short distance cross section for quark-antiquark production can be calculated perturbatively. However, the subsequent evolution of the quark-antiquark pair into a bound quarkonium state is governed by nonperturbative dynamics. The associated long-distance matrix elements (LDMEs) must therefore be determined from fits to data. A further complication arises in the low $p_{T}$ region, $p_{T}\lesssim M_{Q}$, where pQCD is reliable and nonperturbative effects may dominate. Fortunately, most of the nonperturbative effects cancel in cross section ratios measured under identical kinematical conditions, allowing for more direct comparison with theory.
The earliest exploratory experiments [70, 71, 32], primarily focused on deviations from unity of the exponent $\alpha$ extracted from data. Measurements at CERN [72, 73] and Fermilab [74, 75] demonstrated that a constant $\alpha$ could not adequately describe the observed $x_{F}$ and $p_{T}$ spectra. Detailed proton-induced measurements on several nuclear targets conducted by the Fermilab E772 [76] and E866 [77] collaborations confirmed the strong variation of $\alpha$ in both ${J}/\psi$ and $\psi^{\prime}$ production.
@@
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Process Target ($A$) Dependence Ref.
@@
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Process Target ($A$) Dependence Ref.
@@
A few years later, the E866 collaboration [80] repeated the E772 studies by significantly extending the kinematic coverage of the data. A substantial variation of $J/\psi$ suppression as a functions of both $x_{F}$ and $p_{T}$ was observed. Moreover, the data showed a distinct suppression patterns for $J/\psi$ and $\psi^{\prime}$, for all positive values of $x_{F}$. This suppression becomes more pronounced at large $x_{F}\gtrsim 0.3$. Complementary high-statistics measurements by the NA50 collaboration at CERN confirmed [43] that the loosely bound $\psi^{\prime}$ is more strongly suppressed than the $J/\psi$. Further investigation using a higher-intensity proton beam showed [82] that the ratio of $\psi^{\prime}$ to $J/\psi$ production significantly decreases with increasing $A$, also indicating stronger $\psi^{\prime}$suppression.
@@
The NA50 collaboration at CERN investigated $J/\psi$ suppression using seven nuclear targets with proton beam energies of 158 and 400 GeV. Their analysis [85] presented as a function of the thickness of the nuclear target traversed by the $J/\psi$, $L$, revealed a significant $\sqrt{s_{NN}}$ dependence that became more pronounced for larger $L$. This energy dependence remained when the data were presented as a function of $x_{2}$.
More recently, heavy flavor production measurements in fixed-target mode with the LHC proton beam at 2.5, 4.0 and 6.5 GeV were reported. Using a dedicated internal gas target (SMOG), the LHCb collaboration was able to extend [87, 86] the fixed-target center-of-mass domain covered to $\sqrt{s_{NN}}=110.4$ GeV. These first and promising measurements are in good agreement with previous data at lower energies.
Subsequent measurements at the LHC have confirmed the suppression patterns observed in fixed-target experiments at lower beam energies. In $p+A$ collisions, the nuclear modification factor $R_{pA}$ for $J/\psi$ exhibits stronger suppression at forward rapidity compared to backward rapidity, indicating a rapidity-dependent nuclear effect. This behavior is consistent across measurements by ALICE [94, 95, 96, 97, 98, 99], LHCb [100, 101, 103, 104, 105, 102], and CMS [106, 107, 108, 109]. Furthermore, the suppression of different quarkonium states varies with the strength of their binding energies. The hierarchy of suppression observed follows the trend $R_{pA}^{\psi({\rm 2S})}<R_{pA}^{J/\psi}<R_{pA}^{\Upsilon({\rm 1S})}$, where the $\psi$(2S) states are more suppressed and the $\Upsilon$(1S) states being least suppressed [110, 111, 112]. States with lower binding energies, like the $\psi$(2S), are thus more easily dissociated. This hierarchy is more pronounced at backward rapidity than at forward rapidity. In addition, the suppression of $\Upsilon$ states follows the hierarchy $R_{pA}^{\Upsilon({\rm 1S})}<R_{pA}^{\Upsilon({\rm 2S})}<R_{pA}^{\Upsilon({\rm 3S})}$ [109]. The relative suppression among the $\Upsilon$ excited states is also more pronounced at backward than at forward rapidity. In summary, the suppression patterns of quarkonium states in $p+A$ collisions at the LHC are rapidity-dependent and reflect the interplay between CNM effects, quarkonium masses and binding energies.

### v2

@@ Abstract
This review examines data from hadron-nucleus collisions, primarily focusing on hard processes like Drell-Yan, heavy flavor and quarkonium production. It highlights observed modifications of particle yields as functions of momentum and rapidity, aiming to clarify the underlying QCD effects on cold nuclear matter. It outlines strategies for future experiments, including the Electron-Ion Collider, to distinguish between these effects. Key questions address the universality of suppression mechanisms and the role of nonperturbative physics, providing a road map for upcoming measurements of hadrons on nuclei.
@@ I Introduction
The root cause of this unsatisfactory situation lies in the multitude of nuclear effects that can influence the nuclear cross sections. The well known EMC effect [1] provided evidence that the nuclear parton distributions functions (nPDFs) in bound nucleons differ from those in free nucleons, modifying the effective parton densities probed in the collision [2, 3, 4, 5, 6, 7, 8, 9]. Most hadron-nucleus experiments primarily probe the leading-twist shadowing region of the target nucleus at low momentum fraction $x$, leading to a reduction of the particle yields relative to an incoherent sum over all the nucleons.
@@
Another effect, relevant at high center-of-mass energies, is due to gluon saturation. Saturation prevents an unlimited increase of the gluon density at low $x$ values, which would otherwise violate unitarity constraints. This saturation has an important impact on particle production, resulting in additional suppression at very forward rapidities. This suppression is similar to the suppression resulting from the purely nonperturbative leading-twist shadowing. Although saturation and shadowing originate from two different QCD mechanisms, differentiating between them remains a difficult task.
@@
In addition, the incoming or outgoing parton may lose some of its energy by medium-induced gluon radiation when traversing the nucleus, known as Landau-Pomeranchuk-Migdal (LPM) energy loss  [10, 11]. Similarly, the incoming and outgoing partons could undergo energy loss, which is summed coherently and is known as Fully Coherent Energy Loss (FCEL)  [12, 13]. A related, but distinct, phenomenon arises from multiple soft scatterings of the parton traversing the nuclear medium. Each consecutive scattering contributes to a slight modification of the transverse momentum of the parton as well as the broadening of the particle production yields as a function of transverse momentum, $p_{T}$ [14, 15, 16, 17, 18, 19]. Since the broadening scales with the number of scatterings, it typically grows linearly with the nuclear size.
@@
Furthermore, heavy mesons produced in the final state are subject to inelastic interactions with the nuclear medium. When their formation length is shorter than the size of the nucleus, the newly formed mesons may be dissociated via scattering with partons in a high-density environment [20, 21]. In addition, both the initially produced $Q\overline{Q}$ pair and the nascent heavy meson can interact with the surrounding nucleons. The corresponding absorption cross section can vary significantly depending on the energy and production mechanism. Moreover, the quarkonium state may undergo further interactions with comoving hadrons or partons, referred to comovers [22, 23]. These interactions, which result in the break-up of the bound state, are affected by the local density of the medium. The effects of collisional dissociation, nuclear absorption, and comover interactions are strongly dependent on the path length of the state through the nuclear medium and all lead to attenuation of the observed heavy meson yields. In a typical high-energy hadron-nucleus experiment, several CNM effects may be present simultaneously. Their relative importance, however, depends strongly on the process considered, the kinematic regime, and the specific observables under scrutiny. The difficulties in quantitatively establishing the relative contributions of these effects adds to the challenge of forming a coherent theoretical framework.
@@
The upcoming Electron-Ion Collider (EIC) [24] and future fixed-target experiments [25, 26] will offer unprecedented opportunities to probe CNM with greater precision and across a broad kinematic range. Leveraging these new experimental capabilities, we seek to narrow the gap between theoretical predictions and experimental observations, paving the way for a unified description of cold nuclear matter effects.
We identify three key questions that future data should address:
@@
This paper is organized in eight sections. The most relevant fixed-target and collider data are reviewed in Sec. II. Section III discusses the present status of nPDFs and gluon saturation at small $x$. Radiative energy loss effects are reviewed in Sec. IV. Final-state interactions including collisional dissociation, absorption, and comovers are discussed in Sec. V. Additional effects due to impact-parameter dependence and intrinsic charm are presented in Sec VI. Section VII explores the experimental opportunities opened by the newest physics facilities and, particularly, by the EIC. A short conclusion is drawn in Sec. VIII.
@@ II Drell-Yan and heavy flavor data on nuclear targets
Hadron-nucleus collisions serve as a powerful tool to probe the properties of cold nuclear matter, offering unique insights into the dynamics of partonic interactions in the nuclear medium. Assuming the validity of collinear factorization, they provide a useful context for extracting and studying nPDFs. Processes such as Drell-Yan and single inclusive hadron production are particularly valuable probes of these effects. Fixed-target experiments, which benefit from a large experimental rapidity, $y$, coverage due to the beam boost, primarily explore nuclear effects at higher values of the parton momentum fraction $x$, $x>10^{-3}$. In contrast, data from the Large Hadron Collider (LHC) extend these studies into the low $x$ regime, even though they cannnot cover the full rapidity range available, reaching $x$ values as small as $\sim 10^{-5}$, where gluon saturation is expected to become significant. The nuclear modification factor, $R_{hA}$, quantifies the impact of the nuclear medium on particle production by comparing cross sections in hadron-nucleus, $h+A$, to hadron-proton, $h+p$, collisions. Studied as a function of rapidity or Feynman $x$ ($x_{F}$) and transverse momentum, $p_{T}$, $R_{hA}$ provides insight into the evolution of nuclear effects across kinematic regimes.
Despite extensive measurements at fixed-target facilities and colliders, the mechanisms driving the nuclear effects remain ill-determined. It does not appear possible to precisely determine the contributions from all proposed cold nuclear matter effects based on only one type of measurement. Gaining a more quantitative insight into these effects requires a combined analysis of various processes. The current status of the main processes used to investigate cold nuclear matter effects—namely Drell-Yan, quarkonium and heavy-flavor production in hadronic collisions, as well as particle production in ultraperipheral collisions—is summarized below.
@@ II.1 Drell-Yan Data
At leading order (LO), the electromagnetic Drell-Yan process describes the production of a virtual photon, $\gamma^{\star}$, via quark-antiquark annihilation and its decay into a dilepton (lepton-antilepton pair), $\gamma^{\star}\to\ell^{+}\ell^{-}$. Along with deep-inelastic scattering (DIS), this process provides complementary information about the partonic structure of both the beam particle and the target proton or nucleus. The Drell-Yan cross section is well understood and can be calculated using fixed-order perturbative methods [27, 28, 29] or employing advanced approaches such as Collins-Soper-Sterman (CSS) resummation and transverse-momentum-dependent (TMD) factorization [30], which accurately describes the low $p_{T}$ portion of the dilepton spectrum. The nuclear dependence of the Drell-Yan process is quantified by the ratio $R_{hA}^{\text{DY}}$, which compares the yields in $h+A$ collisions to those in $h+p$ or, occasionally, in $h+{\rm d}$ collisions.
@@
Most of the Drell-Yan data come from fixed-target experiments, primarily performed at CERN and Fermilab, spanning the center-of-mass energy range $15\leq\sqrt{s}\leq 38.7$ GeV. The earliest experiments [31, 32, 33], which used proton and pion beams at 225 GeV and measured only total cross sections, showed no evidence of nuclear effects. This conclusion was based on comparisons of $p+p$ and $p+A$ cross sections using the relation $\sigma_{pA}=A^{\alpha}\sigma_{pp}$, where the exponent $\alpha$ was found to be consistent with unity. More detailed comparisons between heavy and light targets were performed in the beginning of the 1980s using pion beams at CERN [34] and proton beams at Fermilab [35]. In spite of the relatively large statistics, both experiments reported the absence of nuclear effects. A clear variation of $R_{hA}^{\text{DY}}$ from unity was observed [36] a few years later by the NA10 experiment at CERN, when comparing the data collected on hydrogen and tungsten targets. [^1] The nuclear modification factor $R_{hA}$ can be related to $\alpha$ on the cross section level by $R_{hA}=\sigma_{pA}/(A\sigma_{pp})=A^{\alpha-1}$. The measured $R_{hA}^{\rm DY}$ as a function of the target parton momentum fraction, $x_{2}$, was found to agree with the recently discovered EMC effect [1]. A pronounced increase of $R_{hA}^{\text{DY}}$ around $p_{T}=2$ GeV (the “Cronin peak” [37]) was also observed [38]. The NA10 results were soon qualitatively confirmed by extensive measurements by the E772 experiment at Fermilab [39]. Taking data with a 800 GeV proton beam, the experiment compared results from a deuterium target with four other nuclear targets (C, Ca, Fe and W) in the $x_{F}$ region between 0.1 and 0.3.222Note that Feynman $x$, $x_{F}$, can be defined as $x_{F}=x_{1}-x_{2}$ the difference between the projectile momentum fraction, $x_{1}$ and the target momentum fraction, $x_{2}$, with a range $-1\leq x_{F}\leq 1$. Another definition is $x_{F}=p_{z}/p_{z}^{\text{max}}$, where $p_{z}$ is the momentum of the dilepton along the beam direction, and $p_{z}^{\text{max}}=\sqrt{s}/2$. The data were found to be consistent with unity, except at low $x_{2}$, where the suppression increased with the mass of the target, $A$, an effect fairly compatible with the leading-twist nuclear shadowing measured in DIS. An increase of $R_{hA}^{\text{DY}}$ for $p_{T}\gtrsim 1$ GeV with $A$ was also clearly observed. A much larger $x_{F}$ region, up to $x_{F}=0.95$, was covered by the E866 collaboration [40] using three targets: Be, Fe and W. These data thus extended to larger values of $x_{1}$ and thus to smaller values of $x_{2}$, providing increased sensitivity to nuclear effects in these two domains. At low $x_{2}$, the experimental Fe/Be and W/Be ratios are compatible with the leading-twist nuclear shadowing observed in DIS while for $x_{1}\gtrsim 0.7$ the observed suppression provided the first evidence for parton energy loss in the projectile in the Drell-Yan process [41, 42, 4].
@@
High statistics Drell-Yan data employing a 450 GeV proton beam was obtained by the NA50 collaboration at CERN [43]. When corrected for nuclear isospin, the resulting cross sections for a number of nuclei from Be to W perfectly scale with the number of nucleon-nucleon collisions, in line with previous data on proton and deuteron targets. More recently, preliminary data [44] from the E906 experiment at Fermilab (SeaQuest) using a proton beam at 120 GeV became available. The measured nuclear modification factor $R_{pA}^{\text{DY}}$ was found to be smaller than unity as a function of $x_{F}$, with a stronger suppression at large $x_{F}\sim 0.8$. The modification appears to be more pronounced for W/C ratio than for Fe/C.
@@
Measurements of the dilepton continuum have been extended by several collider experiments. The PHENIX collaboration at RHIC [45] reported data for $p+p$ and $p+\mathrm{Au}$ collisions at $\sqrt{s_{NN}}=200$ GeV, focusing on the dilepton mass region below 10 GeV and $p_{T}\lesssim M$. A slight enhancement in the $p_{T}$ spectra was observed at forward rapidity around $p_{T}=2$ GeV. Due to the low statistics, these data do not exhibit any clear nuclear effects.
@@
The study of the Drell-Yan continuum is complemented by $Z$ and $W^{\pm}$ vector boson measurements made by the LHC experiments, see recent results from $p+{\rm Pb}$ collisions by CMS [47, 48, 49, 46], ALICE [50, 51] and LHCb [52, 53] at 5.02 and 8.16 TeV. Within uncertainties, the ALICE data show only a small difference between the ${p+p}$ and $p+A$ cross sections. In contrast, the LHCb data indicate a slight suppression of the nuclear cross section in the forward region and an enhancement in the backward region compared to data in ${p+p}$ collisions. Both data sets are compatible with the nPDF calculations. This difference may partly arise from the distinct rapidity coverages: ALICE data probe $-4.46<y<-2.96$ and $3.532.03<y<$ 3.53, while LHCb data access slightly different rapidity regions with $-4.0<y<-2.5$ and $4.01.5<y<$ 4.0. The different kinematic ranges could influence the observed nuclear effects.
The Drell-Yan, $Z$ and $W^{\pm}$ vector boson data from the fixed-target and collider experiments discussed above are summarized in Table 1.
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Final State Target ($A$) Observable Ref.
@@
Table 1: List of data sets for Drell-Yan (DY) and $Z$, $W^{\pm}$ vector bosons production on nuclear targets.
@@ II.2 Quarkonium Data
Quarkonium production involves the formation of a heavy quark-antiquark bound state through hadronization, $h+A\to[Q\bar{Q}]+X$. Several approaches have been used to describe this process, including the Color Evaporation Model (CEM) [57, 58, 59, 60], the Improved Color Evaporation Model (ICEM) [61], the Color Singlet Model (CSM) [62], and Non-Relativistic Quantum Chromodynamics (NRQCD) [63]. The CEM and CSM have not be rigorously applied beyond hadroproduction. The ICEM has been successfully been applied to production and polarization of $J/\psi$ and $\Upsilon$ in $p+p$ and $J/\psi$ in $e+p$ collisions [64, 65, 66, 67, 68, 69]. Work is ongoing to apply it to other quarkonium states and production mechanisms. Although theoretically sound and phenomenologically successful, the NRQCD approach appears incomplete, as outlined in the next paragraph.
In the NRQCD framework, the short distance cross section for quark-antiquark production can be calculated perturbatively. However, the subsequent evolution of the quark-antiquark pair into a bound quarkonium state is governed by nonperturbative dynamics. The associated long-distance matrix elements (LDMEs) must therefore be determined from fits to data. A further complication arises in the low $p_{T}$ region, $p_{T}\lesssim M_{Q}$, where pQCD is less reliable and nonperturbative effects may dominate. Fortunately, most of the nonperturbative effects cancel in cross section ratios measured under identical kinematical conditions, allowing for more direct comparison with theory.
The earliest exploratory experiments [70, 71, 32], primarily focused on deviations of the exponent $\alpha$, extracted from the data, from unity. Measurements at CERN [72, 73] and Fermilab [74, 75] demonstrated that a constant $\alpha$ could not adequately describe the observed $x_{F}$ and $p_{T}$ spectra. Detailed proton-induced measurements on several nuclear targets conducted by the Fermilab E772 [76] and E866 [77] collaborations confirmed the strong variation of $\alpha$ in both ${J}/\psi$ and $\psi^{\prime}$ production.
@@
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Final State Target ($A$) Observable Ref.
@@
Exp. Beam $\sqrt{s_{NN}}$ (GeV) Final State Target ($A$) Observable Ref.
@@
A few years later, the E866 collaboration [80] repeated the E772 studies by significantly extending the kinematic coverage of the data. A substantial variation of $J/\psi$ suppression as a functions of both $x_{F}$ and $p_{T}$ was observed. Moreover, the data showed distinct suppression patterns for $J/\psi$ and $\psi^{\prime}$, for all positive values of $x_{F}$. This suppression becomes more pronounced at $x_{F}\gtrsim 0.3$. Complementary high-statistics measurements by the NA50 collaboration at CERN confirmed [43] that the loosely bound $\psi^{\prime}$ is more strongly suppressed than the $J/\psi$. Further investigation using a higher-intensity proton beam showed [82] that the ratio of $\psi^{\prime}$ to $J/\psi$ production significantly decreases with increasing $A$, also indicating stronger $\psi^{\prime}$suppression.
@@
The NA50 collaboration at CERN investigated $J/\psi$ suppression using seven nuclear targets 158 and 450 GeV proton beams. Their analysis [85], presented as a function of the thickness of the nuclear target traversed by the $J/\psi$, $L$, revealed a significant $\sqrt{s_{NN}}$ dependence that became more pronounced for larger $L$. This energy dependence remained when the data were presented as a function of $x_{2}$.
More recently, heavy flavor production measurements in fixed-target mode with the LHC proton beams of 2.5, 4.0 and 6.5 GeV ($\sqrt{s}/2$ in collider mode) were reported. Using a dedicated internal gas target (SMOG), the LHCb collaboration was able to fill gaps in the fixed-target energy reach from $\sqrt{s_{NN}}=69$ to 110.4 GeV [87, 86]. These first, promising, measurements are in good agreement with previous data at lower energies.
Subsequent measurements at the LHC have confirmed the suppression patterns observed in fixed-target experiments at lower beam energies. In $p+A$ collisions, the nuclear modification factor $R_{pA}$ for $J/\psi$ exhibits stronger suppression at forward rapidity compared to backward rapidity, indicating a rapidity-dependent nuclear effect. This behavior is consistent across measurements by ALICE [94, 95, 96, 97, 98, 99], LHCb [100, 101, 103, 104, 105, 102], and CMS [106, 107, 108, 109]. Furthermore, the suppression of different quarkonium states varies with the strength of their binding energies. The hierarchy of suppression observed follows the trend $R_{pA}^{\psi({\rm 2S})}<R_{pA}^{J/\psi}<R_{pA}^{\Upsilon({\rm 1S})}$, where the $\psi$(2S) states are more suppressed and the $\Upsilon$(1S) states being least suppressed [110, 111, 112]. States with lower binding energies, like the $\psi$(2S), are thus more easily dissociated. This hierarchy is more pronounced at backward rapidity than at forward rapidity. In addition, the suppression of $\Upsilon$ states follows the hierarchy $R_{pA}^{\Upsilon({\rm 3S})}<R_{pA}^{\Upsilon({\rm 2S})}<R_{pA}^{\Upsilon({\rm 1S})}$ [109]. The relative suppression among the $\Upsilon$ excited states is also more pronounced at backward than at forward rapidity. In summary, the suppression patterns of quarkonium states in $p+A$ collisions at the LHC are rapidity-dependent and reflect the interplay between CNM effects, quarkonium masses and binding energies.